OpenStudy (anonymous):

If a circle has a radius of 8 cm. what is the length, in centimeters, of the arc that is intercepted by a central angle of 60 degrees?

OpenStudy (anonymous):

A central angle is measured by its intercepted arc. Let's denote the length of the intercepted arc with s, and the length of the radius r. So, s = 6 cm and r = 30 cm. When a central angle intercepts an arc whose length measure equals the length measure of the radius of the circle, this central angle has a measure 1 radian. To find the angle in our problem we use the following relationship: measure of an angle in radians = (length of the intercepted arc)/(length of the radius) measure of our angle = s/r = 6/30 = 1/5 radians. Now, we need to convert this measure angle in radians to degrees. Since pi radians = 180 degrees, then 1 radians = 180/pi degrees, so: 1/5 radians = (1/5)(180/pi) degrees = 36/pi degrees, or approximate to 11.5 degrees.

OpenStudy (anonymous):

did u get it

OpenStudy (anonymous):

the list of possible answers given are in terms of pi 4pi/3 8pi/3 4pi 8pi 10pi